If X is the number obtained when a die is thrown, show that Tchebycheff's inequality gives P(|X-3.5|>2.5)<0.47. Compare this with actual probability


1
Expert's answer
2021-02-03T05:04:50-0500

We compute the mean ("\mu") and the standard deviation ("\sigma") for "X". We get:

"\mu=E[X]=\frac16(1+2+3+4+5+6)=3.5"

"\sigma=E[X^2]-(E[X])^2=\frac16(1^2+2^2+3^2+4^2+5^2+6^2)-(E[X])^2=\frac{91}6-\frac{49}4=\frac{35}{12}"

We remind that Chebyshev inequality states that "P(|X-\mu|\geq k\sigma)\leq\frac{1}{k^2}". Another version of Chebyshev inequality is: "P(|X-\mu|\geq k)\leq\frac{\sigma^2}{k^2}"

We consider the second version of the inequality. It can be obtained from Markov inequality: "P(X\geq c)\leq\frac{E[X]}{c}". We apply Markov inequality to random variable: "(X-\mu)^2" and receive: "P(|X-\mu|\geq k)=P((X-\mu)^2\geq k^2)\leq\frac{E[(X-\mu)^2]}{k^2}=\frac{E[X^2]-2\mu E[X]+\mu^2}{k^2}=\frac{\sigma^2}{k^2}"

In a similar way one can obtain the first version.

In both cases "k" is a strictly positive real number.

We set "\mu=3.5" and "\sigma=\frac{35}{12}\approx2.92" and receive: "P(|X-3.5|\geq\,2.92\,k)\leq\frac{1}{k^2}" and

"P(|X-3.5|\geq k)\leq\frac{(2.92)^2}{k^2}". We set "k\approx0.856" in the first inequality and "k=2.5" in the second one. We receive: "P(|X-3.5|\geq\,2.5)\leq1.168" from the first inequality and "P(|X-3.5|\geq k)\leq1.364" from the second one. Now we shall calculate the probabilities: "P(|X-3.5|\geq\,2.5)" and "P(|X-3.5|>\,2.5)". We receive:

"P(|X-3.5|\geq\,2.5)=P((X=5)\lor(X=6))=\frac{1}{3}" and

"P(|X-3.5|>\,2.5)=P(X=6)=\frac{1}{6}".

As it can be observed, two versions of Chebyshev inequality provide estimates 1.168 and 1.364, whereas direct calculations yield "\frac{1}{3}" (or "\frac{1}{6}" if the inequality is strict).